# Social Norms in Cinema: A Cross-Cultural Analysis of Shame, Pride and Prejudice

### v4

@@ Abstract
Shame and pride are social emotions expressed across cultures to motivate and regulate people’s thoughts, feelings, and behaviors. In this paper, we introduce the first cross-cultural dataset of over $10k$ shame/pride-related expressions, with underlying social expectations from  5.4K Bollywood and Hollywood movies. We examine how and why shame and pride are expressed across cultures using a blend of psychology-informed language analysis combined with large language models. We find significant cross-cultural differences in shame and pride expression aligning with known cultural tendencies of the USA and India -- e.g., in Hollywood, shame-expressions predominantly discuss self whereas Bollywood discusses shame toward others. Pride in Hollywood is individualistic with more self-referential singular pronouns such as I and my whereas in Bollywood, pride is collective with higher use of self-referential plural pronouns such as we and our. Lastly, women are more sanctioned across cultures and for violating similar social expectations e.g. promiscuity.
@@ 1 Introduction
Shame and pride are social emotions expressed across cultures to motivate and regulate people’s thoughts, feelings, behaviors, achievements, and social adaptations Fischer and Tangney (1995); Goetz and Keltner (2007); Fessler (2007); Schaumberg and Skowronek (2022). The expressions and appraisals of shame and pride vary across cultures and is a research question of great significance to understand society and self-regulation Wong and Tsai (2007); Lewis et al. (2010); Wegge et al. (2022). The overarching question we ask in this paper is, how and why shame and pride are expressed across cultures. Knowing cross-cultural differences in shame/pride-related expressions will help us measure psychological constructs such as shame proneness Tangney et al. (1992); Cohen et al. (2011) and self-esteem Pyszczynski et al. (2004) reliably from language -- enabling effective human-computer interactions in domains such as AI-driven psychotherapy Resendiz and Klinger (2023); Demszky et al. (2020). Examining behaviors evoking shame and pride will reveal culture-specific societal expectations and attitudes (See Table 1) -- paramount to building culturally competent AI models for diverse users Talat et al. (2021); Atari et al. (2023).
Look, Alok is in love with me and I love him too.
Shameless one! Have you lost all your shame?
What is there to be ashamed in this? Don’t sisters love their brothers in this vicinity?
$\Rightarrow$ Incestuous relationship evoke shame.
@@
For our study, we select two top-grossing movie industries, i.e., Hollywood and Bollywood  [^1] https://nairametrics.com/2024/08/14/top10-fastest-growing-film-industries-in-the-world/. Hollywood primarily depicts social situations from the USA, an individualist society that values competency and autonomy Triandis (1989, 1988). Bollywood depicts India, a collectivist society where one’s sense of self is interwoven with community beliefs. The cultural dichotomy between India and the USA (i.e., collectivism vs. individualism) thus presents a rich ground for understanding variations in their beliefs and values. Our approach blends psychology-informed language analysis with state-of-the-art Large Language Model (LLM) to delineate interpretable psychosocial constructs such as self-focus, morality, etc. in shame/pride-related expressions and extract implicit social expectations behind them.
This paper makes the following research contributions: (a) We build a cross-cultural dataset of over 10k shame/pride-related dialogues with underlying social expectations  [^2] Link will be provided upon acceptance, (b) We demonstrate significant cross-cultural variations in how and why shame and pride are expressed in movies, and (c) We illustrate that women are more subjected to social sanctions than men across cultures and for violating similar social expectations.
To the best of our knowledge, this is the first empirical study analyzing the psycho-social constructs underlying shame/pride-related expressions across cultures and the social expectations behind them.
@@ 2 Movies Subtitles Corpora
Movies provide culture-specific life-like depictions of social situations Adkins and Castle (2014); Kubrak (2020). The natural conversation style between characters in movies can reveal the social power dynamics (e.g., boss-employee, father-daughter) and gender roles. Another potential alternative is social media posts however, they are (a) skewed to social situations specific to young demographic groups from Western countries, and (b) are second-hand reports that do not reveal fully the social dynamics of how shame/pride are expressed in natural conversations, for instance, between a boss and an employee or among family members at a Christmas table.
We collected English subtitles for 5,435 Hollywood and Bollywood movies that were released post-1990 by auto-crawling websites that host or link movie subtitles (See Table 2 for data distribution). The subtitles are professionally done translations to preserve the emotions and intended message. Bollywood movie subtitles were collected from www.Bollynook.com and Hollywood movie subtitles were from the publicly available Kaggle movie subtitles dataset [^3] https://www.kaggle.com/datasets/adiamaan/movie-subtitle-dataset. The year of release for movies was verified by either parsing subtitle file names having a release year or checking Wikipedia entries. The year mapping was performed to ensure a similar period for collected movies.
@@
Table 2: Data Distribution. #shame indicates the number of dialogues with the word "shame" or its derivative form (e.g., ashamed, shameless - See Table A1). Similarly, #pride indicates the number of dialogues with the word pride or its derivative form (e.g., proud). Control is the group of dialogues without words shame, pride, and their derivatives.
@@ 2.1 Extracting Shame and Pride related Expressions
The expressions of shame and pride could be explicit (See Table 1) or implicit (e.g. characters’ lowered shoulders, avoiding eye contact, etc.). We extracted explicit mentions of shame and pride to learn about occurrences of shame and pride as determined by the characters. We adopted a keyword search-based approach (See Appendix A1) to identify dialogues along with the previous and the next two lines for situational context (See Table 1). In the case of short dialogues, which could be the case due to monosyllabic responses in spoken conversations, an extra previous and next line was appended to context. One author manually checked the entire dataset of shame and pride-related dialogues to filter out dialogues with phrases such as "what a shame", "it’s a shame", "proudly presents".
Detecting implicit shame/pride-related expressions in textual discourse is an innately complex task. Computational systems for detecting social emotions such as shame and guilt underperform significantly compared to basic emotions such as joy Demszky et al. (2020); Resendiz and Klinger (2023) due to internal self-evaluation associated with social emotions. We tested the GPT-4 language model Achiam et al. (2023) to detect implicit expressions of shame in given movie dialogues without explicit mentions of shame and pride. We recruited two human annotators to determine the agreement with GPT-4 generated labels (See Appendix A.2 for details). On a randomly sampled 100 situations each from H/Bollywood, the model returned 20 samples to be expressing shame in Bollywood and 5 samples in Hollywood. We found poor agreement between human annotators (4 out of 20 samples in Bollywood and 1 out of 5 samples in Hollywood where both annotators agreed with GPT-4). Human annotators across cultures tend to cross-label shame and guilt Troiano et al. (2019) and follow-up discussions with annotators revealed similar tendencies. One such example from Hollywood is provided below:
‘‘about feeding the poor, I’ve never done any of that. God tells us to love everybody. I’ve hated people… my family, my family, ’’
@@
For our analysis, we created four sets of dialogues: (a) shame-related dialogues in Bollywood, (b) shame-related dialogues in Hollywood, (c) pride-related dialogues in Bollywood and (d) pride-related dialogues in Hollywood. Cross-cultural differences also exist in speaking styles across cultures and we thus formed a control set of dialogues unrelated to shame and pride for both movie industries to remove the variations in language markers owing to culture-specific speaking styles (See Table 2 for dialogue distribution).
@@
Hollywood: LIWC Category (Top-5 words) $r$ Bollywood: LIWC Category (Top-5 words) $r$
Moral (wrong, excuse, decent, honest, duty) 0.110 $2^{nd}$ person pronouns (you, your, you’re, yourself, you’ve) 0.161
$1^{st}$ person sing. pronouns (i, me, i’m, my, i’ll) 0.102 Social References (you, your, he, her, him ) 0.148
Sadness (crying, cry, sob, lonely, sad) 0.096 Moral (wrong, innocent, duty, decent, excuse) 0.102
$2^{nd}$ person pronouns (you, your, you’re, yourself, you’ve) 0.080 Female ( her, she, girl, she’s, mom) 0.060
Past focus (was, did, were, been, didn’t ) 0.069 Family (son, married, uncle, dad, mom) 0.055
All or none (no, all, nothing, never, yes) 0.055 Sexual (chaste, lust, sex, sexy, pimp) 0.051
Family (son, dad, baby, mom, mama) 0.054 Conflict (kill, killed, accusing, killing, cruel) 0.024
Social References (you, your, we, he, you’re) 0.050 Swear words (hell, bloody, idiot, damn, ass) 0.024
Anxiety (fear, afraid, worry, terrified, scared) 0.042 Anger (mad, angry, hate, cruel, argue) 0.022
Illness (sick, pain, pains, flu, sickly) 0.038 Present focus (is, are, don’t, i’m, aren’t) 0.021
Table 3: Pearson r for LIWC categories significantly correlated ($p<0.05$) with shame for Hollywood and Bollywood after Benjamini-Hochberg p-value correction (See Appendix for confidence intervals and p values). Note, $1^{st}$ person sing pronouns are strongly correlated with Hollywood-shame whereas the correlation with $2^{nd}$ person pronouns and social references are up to 3 times stronger in Bollywood compared to Hollywood. See Table A7 and A9 for the complete set of correlations.
@@ 3 Approach
Fig. 1 illustrates our study design comprising a blend of two key approaches (a) a psychosocial vocabulary approach to measure how expressions of shame and pride differ and (b) prompting LLM approach to extract reasons behind them and how they differ between India and the U.S.A.
3.1 Vocabulary Approach: LIWC
Linguistic Inquiry of Word Count (LIWC) Boyd et al. (2022) is a corpus analysis tool widely used in psychology to identify psychological categories of words such as "self-focus" (e.g., I, my, mine) and "other-focus" (e.g., you, yours, theirs) which are of great significance for examining social behaviors such as self-regulation and conformity.
To understand cross-cultural linguistic variations in the manifestation of shame and pride, we computed the normalized distribution of psycho-social categories from the dialogues and examined their correlation with shame and pride compared to the control set. The search keywords used for building social emotions corpus (See Table A1) were removed from the LIWC dictionary to prevent overestimation of shame- (e.g., negative emotion) and pride-related categories (e.g., achievement).
3.2 Pretrained LLM: GPT-4 Chat
@@
Hollywood: LIWC Category (Top-5 words) $r$ Bollywood: LIWC Category (Top-5 words) $r$
Social References (you, your, we, he, you’re) 0.131 Drive (we, our, us, sir, work) 0.122
Family (son, dad, baby, mom, mama) 0.121 Achieve (work, better, win, best, try) 0.098
$1^{st}$ person sing. pronouns (i, i’m, me, my, i’ll) 0.095 Power (sir, own, respect, kill, power) 0.091
$2^{nd}$ person pronouns (you, your, you’re, you’ve, yourself) 0.091 Social References (you, your, he, we, our) 0.085
Male References (he, his, him, man, son) 0.084 Moral (wrong, duty, brave, arrogant, useless) 0.075
Drive (we, our, us, we’re, dad) 0.076 Reward (win, won, glory, success, successful) 0.071
Ethnicity (American, Irish, Chinese, German, Christian) 0.052 Political (nation, army, sultan, president, dynasty) 0.064
Reward (win, won, winner, successful, earned) 0.048 $1^{st}$ person pl. pronouns (we, our, us, we’ll, let’s) 0.060
Achieve (work, better, best, trying, try) 0.045 Family (son, papa, married, dad, uncle) 0.059
Power (sir, own, war, strong, mighty) 0.044 Ethnicity (Indian, Indians, British, Hindi, caste) 0.054
Table 4: Pearson r for LIWC categories significantly correlated ($p<0.05$) with pride for Hollywood and Bollywood after Benjamini-Hochberg p-value correction (See Appendix for confidence intervals and p values). Note the contrast, Achievement-related pride in Bollywood, Self-centered pride vs We-centered pride, Family in Hollywood vs Family in Bollywood See Tables A8 and A10 for all significantly correlated categories.
@@
4 Results
4.1 How is shame/pride expressed across cultures?
Shame is associated with negative emotions, power, and morality in both movie industries (See Table 3), however, significant cultural nuances exist in its manifestation. In Hollywood, shame is more I-focused with remorse (e.g., sadness, anxiety) and past-focused language, whereas in Bollywood, it is more You-focused with frequent social references and present-focused language. In Bollywood, psychosocial categories associated with shame reveal a larger social context (note the contrast in family i.e., nuclear family in Hollywood vs extended family in Bollywood) and social situations (i.e., anger, swear, conflict). Categories female and sexual are exclusively correlated with shame in Bollywood potentially indicating an honor system in collectivist societies Caffaro et al. (2014) whereas it is illness that is uniquely correlated with shame in Hollywood indicating social sanctions around incapability.
A similar dichotomy is observed for pride (See Table 4). Pride in Hollywood is more centered toward family (note, family in Bollywood is male-centric and comprises extended family (such as uncle)) and self-centered (e.g., I am proud of __) whereas it is achievement/power focused and "We"-centered in Bollywood movies (with othering (us vs they)- See Table A8). Male references are more strongly associated with pride than females in Hollywood whereas the female category has an insignificant ($p>0.05$) correlation with pride in Bollywood. Pride is exclusively positive (see POSEMO in Table A10) in Hollywood whereas in Bollywood, we suppose that it relates to honor more (e.g., protecting family’s pride, bringing pride to family) and is not always used in a positive context.
Cross-Cultural Shame/Pride Dataset
For the Bollywood set, GPT-4 chat predicted reasons (also implicit social expectations) for 5321 (98.4%) shame-related dialogues out of 5409, and 2237 (74.6%) pride-related dialogues out of 2999. For the Hollywood set, GPT-4 chat predicted a reason for 1156 (94.6%) shame-related dialogues out of 1221 and 1731 (61.7%) pride-related dialogues out of 2805. Upon manual analysis, we found that pride is also used to express affection, specifically toward close family members such as sons - therefore, a lower number of reasons is associated with pride compared to shame. GPT-4 chat predicted a total of 10,445 social expectations (See Table 5 for distribution). Prompting "reasons" behind shame and pride allowed us to capture high specificity in cultural norms (See Tables 6 for shame-related expectations and A6 for pride-related expectations).
4.2.1 Manual Evaluation
Two volunteers manually verified the predicted gender for "the person experiencing shame/pride" and the predicted "reason" in a randomly sampled set of $100$ dialogues each from Hollywood and Bollywood. The annotator who labeled the Bollywood set is an Indian, aware of social roles and expectations in Indian society. Likewise, the annotator for the Hollywood set is an American with a nuanced understanding of the social expectations of the U.S.A.

### v5

@@ Abstract
Shame and pride are social emotions expressed across cultures to motivate and regulate people’s thoughts, feelings, and behaviors. In this paper, we introduce the first cross-cultural dataset of over $10k$ shame/pride-related expressions, with underlying social expectations from  5.4K Bollywood and Hollywood movies. We examine how and why shame and pride are expressed across cultures using a blend of psychology-informed language analysis combined with large language models. We find significant cross-cultural differences in shame and pride expression aligning with known cultural tendencies of the USA and India – e.g., in Hollywood, shame-expressions predominantly discuss self whereas shame is expressed toward others in Bollywood. Women are more sanctioned across cultures and for violating similar social expectations.
@@
Sunny Rai, Khushang Jilesh Zaveri, Shreya Havaldar, Soumna Nema, Lyle H. Ungar & Sharath Chandra Guntuku University of Pennsylvania, Johns Hopkins University, Mahindra University {sunnyrai, sharathg}@seas.upenn.edu
@@ 1 Introduction
Shame and pride are social emotions expressed across cultures to motivate and regulate people’s thoughts, feelings, behaviors, achievements, and social adaptations Fischer and Tangney (1995); Goetz and Keltner (2007); Fessler (2007); Schaumberg and Skowronek (2022). The expressions and appraisals of shame and pride vary across cultures and is a research question of great significance to understand society and self-regulation Wong and Tsai (2007); Lewis et al. (2010); Wegge et al. (2022). The overarching question we ask in this paper is, how and why shame and pride are expressed across cultures. Knowing cross-cultural differences in shame/pride-related expressions will help us measure psychological constructs such as shame proneness Tangney et al. (1992); Cohen et al. (2011) and self-esteem Pyszczynski et al. (2004) reliably from language – enabling effective human-computer interactions in domains such as AI-driven psychotherapy Resendiz and Klinger (2023); Demszky et al. (2020). Examining behaviors evoking shame and pride will reveal culture-specific societal expectations and attitudes (See Table 1) – paramount to building culturally competent AI models for diverse users Talat et al. (2021); Atari et al. (2023).
And should we bow before others begging....them to marry our daughters? This shall not happen. Neither will the girls be alive here nor shall....we be ashamed of ourselves. You cannot kill the life which God has given. I won’t let you commit the sin.
$\Rightarrow$ Not able to marry off their daughters evokes shame.
@@
For our study, we select two top movie industries, i.e., Hollywood and Bollywood [^1] https://en.wikipedia.org/wiki/Film_industry. Hollywood primarily depicts social situations from the USA, an individualist society that values competency and autonomy Triandis (1989, 1988). Bollywood depicts India, a collectivist society where one’s sense of self is interwoven with community beliefs. The cultural dichotomy between India and the USA (i.e., collectivism vs. individualism) thus presents a rich ground for understanding variations in their beliefs and values. Our approach blends psychology-informed language analysis with state-of-the-art Large Language Model (LLM) to delineate interpretable psychosocial constructs such as self-focus and morality in shame/pride-related expressions and extract implicit social expectations behind them.
This paper makes the following research contributions: (a) We develop a cross-cultural dataset of over 10k shame/pride-related dialogues with underlying social expectations [^2] https://github.com/Khushangz/Cross-Cultural-Social-Norms-Dataset/, (b) We demonstrate cross-cultural variations in how and why shame and pride are expressed in movies, and (c) We illustrate that women are more subjected to social sanctions than men across cultures and for violating similar social expectations.
To the best of our knowledge, this is the first empirical study analyzing the psychosocial constructs underlying shame/pride-related expressions across cultures and the social expectations behind them.
@@ 2 Movies Subtitles Corpora
Movies provide culture-specific life-like depictions of social situations Adkins and Castle (2014); Kubrak (2020). The natural conversation style between characters in movies can reveal the social power dynamics (e.g., boss-employee, father-daughter) and gender roles. Another potential alternative is social media posts however, they are (a) skewed to social situations specific to young demographic groups from Western countries, and (b) are second-hand reports that do not reveal fully the social dynamics of how shame/pride are expressed in natural conversations, for instance, between a boss and an employee.
We collected English subtitles for 5,435 Hollywood and Bollywood movies that were released post-1990 by auto-crawling websites that host or link movie subtitles (See Table 2 for data distribution). The subtitles are professionally done translations to preserve the emotions and intended message. Bollywood movie subtitles were collected from www.Bollynook.com and Hollywood movie subtitles were from the publicly available Kaggle movie subtitles dataset [^3] https://www.kaggle.com/datasets/adiamaan/movie-subtitle-dataset. The year mapping was performed to ensure a similar period for collected movies. The year of release for movies was verified by either parsing subtitle file names having a release year or probing Wikipedia entries.
@@
Table 2: Data Distribution. #shame indicates the number of dialogues with the word "shame" or its derivative form (e.g., ashamed, shameless - See Table A1). Similarly, #pride indicates the number of dialogues with the word pride or its derivative form (e.g., proud). Control comprises dialogues without words shame, pride, or their derivatives.
@@ 2.1 Extracting Shame and Pride related Expressions
The expressions of shame and pride could be explicit (See Table 1) or implicit (e.g. characters’ lowered shoulders, avoiding eye contact, etc.). We extracted explicit mentions of shame and pride to learn about occurrences of shame and pride as determined by the characters. We adopted a keyword search-based approach (See Appendix A1) to identify dialogues along with the previous and the next two lines for situational context (See Table 1). For short dialogues such as monosyllabic responses in spoken conversations, we appended an extra previous and next line to context. One author manually checked the entire dataset of shame and pride-related dialogues to filter out dialogues with conventional phrases such as "what a shame", "it’s a shame", "proudly presents".
Detecting implicit shame/pride-related expressions in textual discourse is a complex task. Computational systems for detecting social emotions such as shame and guilt underperform significantly compared to basic emotions such as joy due to internal self-evaluation associated with social emotions Demszky et al. (2020); Resendiz and Klinger (2023). We used GPT-4 language model Achiam et al. (2023) to detect implicit expressions of shame in given movie dialogues. We recruited two human annotators to determine the agreement with GPT-4 generated labels (See Appendix A.2 for details). On a randomly sampled 100 situations each from H/Bollywood, the model labeled 20 samples to be expressing shame in Bollywood and 5 samples in Hollywood. However, the agreement between human annotators was low (4 out of 20 samples in Bollywood and 1 out of 5 samples in Hollywood where both annotators agreed with GPT-4). Human annotators across cultures tend to cross-label shame and guilt Troiano et al. (2019) and follow-up discussions with annotators revealed similar tendencies. One such example from Hollywood is provided below:
about feeding the poor, I’ve never done any of that. God tells us to love everybody. I’ve hated people… my family, my family,
@@
We created four sets of dialogues: (a) shame-related dialogues in Bollywood, (b) shame-related dialogues in Hollywood, (c) pride-related dialogues in Bollywood and (d) pride-related dialogues in Hollywood. Cross-cultural differences also exist in speaking styles across cultures and we thus formed a control set of dialogues unrelated to shame and pride for both movie industries to remove the variations in language markers owing to culture-specific speaking styles (See Table 2 for dialogue distribution).
@@ 3 Approach
Fig. 1 illustrates our study design comprising two key approaches (a) using a psychosocial vocabulary approach to measure how expressions of shame and pride differ and (b) prompting LLM approach to extract reasons behind them and how they differ between India and the U.S.A.
3.1 Vocabulary Approach
Linguistic Inquiry of Word Count (LIWC) Boyd et al. (2022) is a corpus analysis tool to identify psychological constructs such as "self-focus" and "morality", and therefore, is widely used for examining social behaviors such as self-regulation and conformity.
To understand cross-cultural linguistic variations in the manifestation of shame and pride, we computed the normalized distribution of psychosocial categories in LIWC from the dialogues and examined their correlation with shame and pride compared to the control set. The search keywords used for building social emotions corpus (See Table A1) were removed from the LIWC dictionary to prevent overestimation of shame- (e.g., negative emotion) and pride-related categories (e.g., achievement).
3.2 Prompting Approach
@@
Figure 2: Pearson r for LIWC categories significantly correlated ($p<0.05$) with shame for Hollywood and Bollywood after Benjamini-Hochberg p-value correction (See Appendix for confidence intervals and p values). Note, $1^{st}$ person sing pronouns are strongly correlated with Hollywood-shame whereas the correlation with $2^{nd}$ person pronoun and social references are up to 3 times stronger in Bollywood compared to Hollywood. See Table A7 and A9 for the complete set of correlations.
Figure 3: Pearson r for LIWC categories significantly correlated ($p<0.05$) with pride for Hollywood and Bollywood after Benjamini-Hochberg p-value correction (See Appendix for confidence intervals and p values). Note the contrast, Achievement-related & We-centered pride in Bollywood vs Social & Self-centered pride in Hollywood. See Table A8 and A10 for the complete set of correlations.
4 Results
4.1 How is shame/pride expressed across cultures?
Shame is associated with negative emotions, power, and morality in both movie industries (See Tables A7 and A9) however, significant cultural nuances exist in its manifestation (See Fig. 2). In Bollywood, shame-related expressions are other-oriented as indicated by 2nd person pronouns and social references. Female and sexual categories are correlated with shame exclusively in Bollywood potentially indicating the honor system in collectivist societies Caffaro et al. (2014). Other psychosocial categories such as swearing and conflict with present focus reveal the role of shame in enforcing conformity. In Hollywood, shame expressions are self-focused with remorse (e.g., sadness, anxiety) and have past-focused language. Illness is uniquely correlated with shame in Hollywood indicating social sanctions around incapability.
A similar dichotomy is observed for pride (See Fig. 3). In Bollywood, pride is achievement/power focused and "We"-centered. Pride is associated exclusively with men in Bollywood. Other psychosocial categories such as moral and politics are more strongly associated with pride. In Hollywood, pride is more self-centered and expressed for family more. Male references are more strongly associated with pride than females in Hollywood whereas the female category has an insignificant ($p>0.05$) correlation with pride in Bollywood. Pride has a positive undertone in Hollywood whereas in Bollywood, we speculate that it relates to honor (e.g., protecting family’s pride, bringing pride to family) and does not have a positive undertone.
Cross-Cultural Shame/Pride Dataset
For the Bollywood set, GPT-4 predicted reasons (also implicit social expectations) for 5321 (98.4%) shame-related dialogues out of 5409, and 2237 (74.6%) pride-related dialogues out of 2999. For the Hollywood set, GPT-4 chat predicted a reason for 1156 (94.6%) shame-related dialogues out of 1221 and 1731 (61.7%) pride-related dialogues out of 2805. Upon manual analysis, we found that pride is also used to express affection, specifically toward close family members such as children. As a result, a lower number of reasons is associated with pride compared to shame. GPT-4 chat predicted a total of 10,445 social expectations (See Table 3 for distribution). Prompting "reasons" behind shame and pride allowed us to capture high specificity in cultural norms (See Tables 4 for shame-related expectations and A6 for pride-related expectations).
4.2.1 Manual Evaluation
Two volunteers manually verified the predicted gender for "the person experiencing shame/pride" and the predicted "reason" for a randomly sampled set of $100$ dialogues each from Hollywood and Bollywood. An Indian annotator aware of social roles and expectations in Indian society labeled the Bollywood set. Likewise, an American annotator labeled the Hollywood set.
@@ 4.2.1 Manual Evaluation
4.2.2 Thematic Analysis
Twenty-four clusters for shame-related reasons and fifteen for pride-related reasons were formed using agglomerative clustering ( See Tables A13 and A14 for clustering parameters). The clusters were manually assigned a theme (as depicted on the Y-axis of Fig. 4 and 5) after analyzing the ten closest samples based on cosine distance from the centroid of the cluster (See Table A11 and A12 for examples in each cluster). We computed the relative association for each cluster with Bollywood and Hollywood using eq. 1 and performed the Barnard-Exact Test Barnard (1947) with the Yates Correction Yates (1934) to test if the possibility of observing norms related to pre-assigned themes is statistically different across movie industries.